OpenStudy (anonymous):

Simplify the expression. 1+i ---- 1-i

terenzreignz (terenzreignz):

When you have a complex number at the denominator, you multiply both the numerator and denominator of the expression by its conjugate...

terenzreignz (terenzreignz):

That said... what is the conjugate of 1-i ?

Parth (parthkohli):

Realizing the denominator. XD

terenzreignz (terenzreignz):

Tips? LOL To get the conjugate of a complex number, a + bi, just change the sign that separates the real and imaginary parts... So... for example... conj ( 4 + 5i ) = 4 - 5i conj ( 3 - 2i) = 3 + 2i and so on

OpenStudy (anonymous):

i still dont understand

terenzreignz (terenzreignz):

Okay, a more direct example... suppose we were to simplify \[\Large \frac{5}{3+4i}\] The denominator is 3 + 4i The conjugate of this denominator is 3 - 4i So we multiply the entire expression by... \[\Large \frac5{3+4i}\times \frac{3-4i}{3-4i}\]resulting into \[\Large = \frac{5\cdot(3-4i)}{(3+4i)(3-4i)}= \frac{5(3-4i)}{3^2 - (4i)^2}=\frac{5(3-4i)}{3^2+4^2}\] \[\Large = \frac{5(3-4i)}{25}=\frac15(3-4i)=\frac34-\frac45i\]

terenzreignz (terenzreignz):

Time's not on my side @Indiadabest I have to go now, I actually just wanted to peek into OS :) If you're still having problems, I'm sure there are plenty of online users that can walk you through this :D For now ---------------------------------------- Terence out